### v1

# Integrated Multimodal Sensing and Communication: Challenges, Technologies, and Architectures

Abstract
The evolution towards 6G networks requires the intelligent integration of communication and sensing capabilities to support diverse and complex applications, such as autonomous driving and immersive services. However, existing integrated sensing and communication (ISAC) systems predominantly rely on single-modal sensors as primary participants, which leads to a limited representation of environmental features and significant performance bottlenecks under the emerging requirements of 6G applications. This limitation motivates a paradigm shift from single-modal to multimodal ISAC. In this article, we first analyze the key challenges in realizing multimodal ISAC, including the fusion of heterogeneous multimodal data, the high communication overhead among distributed sensors, and the design of efficient and scalable system architectures. We then introduce several enabling technologies, such as large AI models, semantic communication, and multi-agent systems, that hold promise for addressing these challenges. To operationalize these technologies, we zoom into three architectural paradigms: fusion-based multimodal ISAC (F-MAC), interaction-based multimodal ISAC (I-MAC), and relay-based multimodal ISAC (R-MAC), each tailored to organize devices and modalities for efficient collaboration in different scenarios. Thereafter, a case study is presented based on the F-MAC scheme, demonstrating that the scheme achieves more comprehensive sensing and improves sensing accuracy by approximately 80% compared to conventional single-modal ISAC systems. Finally, we discuss several open issues to be addressed in the future.
Index Terms:
Integrated multimodal sensing and communications; agent AI; semantic communication; large AI model
I Introduction
I-A Background
Integrated sensing and communication (ISAC), as an emerging paradigm, unifies the functions of wireless sensing and communication within a shared hardware and spectral framework [1]. By leveraging common radio frequency (RF) signals for both tasks, ISAC enables efficient resource utilization, reduces hardware redundancy, and enhances spectral efficiency, making it a promising solution for future intelligent wireless systems. However, traditional ISAC approaches predominantly rely on RF signals as the sole modality for sensing, which limits the system’s capability in complex environments [2]. RF signals provide a single-modality perspective, offering limited information, such as range, velocity, and coarse spatial features, which is insufficient for capturing fine-grained semantic or contextual details of complex environments. Additionally, RF-based sensing often struggles to adapt to rapidly changing or highly dynamic environments, as its performance may degrade significantly due to variations in multipath propagation, interference, and occlusions. These limitations promote the paradigm shift from single-modal to multimodal ISAC.
Different types of sensors operate on fundamentally distinct principles and collect sensory data with varying properties. Examples of these sensing techniques include radar, light detection and ranging (LiDAR), red-green-blue-depth (RGB-D) cameras, and the global positioning system (GPS), which are collectively referred to as multimodal sensing. By leveraging the complementary strengths of different modalities, multimodal ISAC holds great promise in enhancing sensing accuracy, robustness, and generalization. For instance, the inclusion of multimodal sensing can accelerate the beam selection process, such as LiDAR-aided beam prediction [3], camera-GPS-aided beam prediction [4], and camera-aided beam prediction [5], where conventional methods require an exhaustive search across all candidate beam pairs. Furthermore, multimodal ISAC systems demonstrate improved adaptability to dynamic scenarios. For example, visual data can compensate for RF degradation under adverse conditions, such as multipath fading or obstruction.
I-B Contributions
Motivated by the limitations of existing single-modal ISAC systems and the unique demands of 6G applications, this article presents several key contributions to promote the design and implementation of efficient and intelligent multimodal ISAC architectures, as follows:
  1. 1.
    Challenge for multimodal ISAC: We conduct a comprehensive analysis of the fundamental challenges in enabling multimodal ISAC. These challenges include (i) the effective fusion of heterogeneous data streams from diverse sensing modalities, (ii) the substantial communication overhead incurred by distributed information exchange among sensors, and (iii) the design of efficient, scalable, and adaptable system architectures capable of supporting real-time sensing and communication under dynamic network conditions.
  2. 2.
    Technologies for multimodal ISAC: We identify and critically examine several promising technologies that serve as enablers for multimodal ISAC, including large AI models (LAMs) [6], semantic communication (SC) [7], and multi-agent systems (MAS) [8]. We discuss their complementary strengths, such as LAMs’ capacity for multimodal fusion, SC’s ability to task-oriented efficient transmission, and MAS’s support for distributed devices, and illustrate how their integration can significantly enhance the semantic understanding and coordination capabilities of multimodal ISAC systems.
  3. 3.
    Architectures for multimodal ISAC: To systematically integrate these technologies, we examine three distinct architectural paradigms tailored to different multimodal collaboration patterns: (i) Fusion-based Multimodal ISAC (F-MAC), which emphasizes centralized semantic fusion; (ii) Interaction-based Multimodal ISAC (I-MAC), which enables direct peer-to-peer semantic interaction among sensors; and (iii) Relay-based Multimodal ISAC (R-MAC), which focuses on the large range and long distance of sensing tasks across modalities. These frameworks provide flexible and scalable design options for deploying ISAC in diverse application environments.
  4. 4.
    Case Validation: We conduct a case study based on the F-MAC architecture to validate the framework. Experimental results demonstrate that the F-MAC scheme achieves enhanced perceptual coverage and improves overall sensing accuracy by approximately 80% compared to conventional single-modal ISAC approaches, thereby showcasing the practical benefits of multimodal integration.
II Challenges for Multimodal ISAC
Despite its considerable potential to enhance environmental perception, the practical realization of multimodal ISAC faces several critical challenges stemming from the heterogeneous nature of sensing modalities and the complexity of real-time communication and computation. These challenges must be carefully addressed to ensure effective system design and deployment across diverse application scenarios.
II-A Heterogeneous Multimodal Fusion
Multimodal ISAC inherently involves the integration of data streams originating from diverse sensing modalities such as RF signals, vision sensors, and LiDAR. These modalities differ substantially in terms of data dimensionality, spatial-temporal resolution, coverage, and semantic abstraction. For instance, aligning 2D visual images with 1D radar waveforms is particularly challenging due to discrepancies in data structure, coordinate systems, and information content. Moreover, inconsistencies in sensing frequency and perception latency further complicate synchronized fusion.
II-B High Communication Overhead
Real-time sensing and decision-making in distributed ISAC systems require frequent and high-volume data exchanges between sensors, edge nodes, and centralized servers. This results in considerable bandwidth consumption and elevated energy expenditure, particularly in wireless or resource-constrained environments. The issue becomes more pronounced in low-dynamic scenarios—such as static nighttime surveillance—where redundant or minimally informative data (e.g., successive identical video frames) continue to be transmitted, thereby wasting transmission resources without improving situational awareness.
II-C Context-Aware Architectural Design
Designing a system architecture that can adapt to diverse operational environments and task requirements remains a fundamental challenge. Rigid deployment modes may offer simplicity in control and resource management but often lack the flexibility needed to accommodate the heterogeneous demands of ISAC applications. For example, while centralized architectures benefit from powerful computing infrastructure, they introduce latency and dependency issues in time-critical or infrastructure-sparse settings. Conversely, purely distributed systems offer autonomy but often suffer from limited coordination and reduced global situational awareness.
Addressing these challenges calls for more intelligent context-aware systems, which is an area where several advanced technologies, such as LAMs, SC, and MAS, offer promising avenues for enabling adaptive multimodal fusion, communication-efficient sensing, and scalable architectural designs.
Figure 1: The illustration of three technologies. (a) Large AI models. (b) Semantic communication. (c) Multi-agent system.
III Technologies for Multimodal ISAC
As shown in Fig. 1, this section introduces several key technologies, detailing how they address the challenges associated with realizing multimodal ISAC as previously discussed.
III-A Large AI Model
LAMs demonstrate exceptional capabilities in processing heterogeneous multimodal data by leveraging their extensive parameter space and rich prior knowledge [6]. As illustrated in Fig. 1(a), inputs from various modalities, such as images, RF signals, and point clouds, are first transformed into token representations through embedding layers. These tokens, combined with learnable positional embeddings, are then fed into a decoder-only transformer architecture. This architecture consists of stacked decoder blocks, each comprising masked multi-head self-attention for autoregressive modeling, followed by feed-forward networks, residual connections, and layer normalization. This token-based representation and fusion strategy enables LAMs to effectively extract and integrate both low-level physical features and high-level semantic information, thereby enhancing performance in tasks such as object recognition and motion estimation.
III-B Semantic Communication
As illustrated in Fig. 1(b), a typical SC system comprises semantic and channel encoders at the transmitter, and corresponding channel and semantic decoders at the receiver [7]. Unlike conventional communication systems that prioritize bit-level fidelity, SC first employs AI-based semantic encoders to extract essential semantics from raw multimodal data, including images, RF signals, and point clouds. Then, only the extracted semantic representations are transmitted, significantly reducing data volume. Finally, at the receiver, AI-based semantic decoders reconstruct task-specific information from the received semantics, such as object categories, positions, etc. This strategy effectively filters out redundant or irrelevant content, enabling efficient operation under stringent bandwidth constraints.
III-C Multi-Agent System
MAS [8] provides the structural foundation for enabling scalable, decentralized intelligence within the multimodal ISAC. In MAS, multiple autonomous agents, either software-based or embodied in hardware, interact with the environment, as well as each other, to collaboratively achieve task objectives. Each agent perceives local conditions, makes decisions, and executes actions, often through cooperative reasoning and intent-aware communication. As shown in Fig. 1(c), the core component of MAS includes four functional modules: a knowledge base for storing domain knowledge, a memory module for accumulating interaction history, a large language model (LLM), such as Deepseek and GPT, for reasoning and decision-making, and external tools for executing actions. During operation, an agent continuously interacts with the environment, stores relevant observations in memory, combines them with contextual knowledge from the knowledge base, and inputs the fused information into the LLM. The LLM then analyzes the situation and invokes appropriate tools (e.g., an AI or generic algorithm) to perform task-specific actions.
Overall, LAM offers robust capabilities for multimodal data processing, analysis, and fusion, as well as for global control and decision-making across heterogeneous sensors. SC ensures high-efficiency and low-overhead data transmission. MAS provides the organizational strategy, endowing individual sensors with intelligence to enable autonomous reasoning and collaborative behavior. While each of these technologies offers distinct advantages, determining how to effectively and appropriately deploy them in varying scenarios remains a critical challenge.
Figure 2: The illustration of the proposed three architectural schemes of the multimodal ISAC. (a) Fusion-based Multimodal ISAC. (b) Interaction-based Multimodal ISAC. (c) Relay-based multimodal ISAC.
IV Architectures for Multi-Modal ISAC
As illustrated in Fig. 2, this section presents three architectural schemes tailored to distinct scenario-specific requirements, in which the identified technologies, LAM, SC, and MAS, are selectively and strategically integrated to enhance both the efficiency and adaptability of the multimodal ISAC.
IV-A Fusion-based Multimodal ISAC
As illustrated in Fig. 2(a), the F-MAC scheme adopts a centralized fusion architecture to integrate semantic information from heterogeneous sensors. Specifically, each sensor is paired with a lightweight edge agent. These edge agents function as semantic encoders, responsible for extracting and compressing task-relevant semantics from local multimodal data. Then, the resulting compact semantic representations are transmitted to a base station for further processing over a wireless channel. At the base station, LAMs serve as a central agent to align and fuse the received multimodal semantics to generate unified representations. The fused features can then be utilized for downstream applications such as global decision-making, collaborative sensing, and environmental modeling. This process is detailed in Fig. 1(a) and Fig. 1(b). A typical application scenario of this solution is intelligent transportation, where cameras, LiDAR sensors, and RF devices are deployed across various intersections [9]. Each sensor, via its corresponding edge agent, extracts semantic information such as vehicle positions, motion trajectories, and Doppler shifts. The compressed semantic features are transmitted to a central LAM located at a traffic control center, where global fusion is performed to achieve an accurate understanding of real-time traffic conditions, thereby enabling intelligent traffic signal coordination and congestion mitigation.
In the F-MAC scheme, the use of a powerful LAM at the central agent enables deep reasoning over globally aggregated data, improving decision accuracy. Meanwhile, SC offloads bandwidth-intensive raw data transmission from the edge and ensures efficiency in bandwidth-limited environments. However, this scheme is heavily reliant on the computational capabilities of the central agent. Consequently, when communication links between edge sensors and the central base station are disrupted, edge agents are unable to operate autonomously. This dependency can lead to reduced system responsiveness and efficiency, particularly in dynamic scenarios with intermittent connectivity.
IV-B Interaction-based Multimodal ISAC
Fig. 2(b) illustrates the I-MAC scheme, which aims to decentralize intelligence across the network fully. In this scheme, each sensor is equipped with a language model-driven edge agent, as depicted in Fig. 1(c). Each edge agent independently processes its multimodal sensory inputs and transmits only essential high-level semantic outcomes, such as alerts, decisions, or event summaries, to neighboring agents. Subsequently, each agent performs local analysis, reasoning, and task-specific decision-making by leveraging the received semantic information, internal memory, and localized knowledge bases. A representative application scenario is industrial inspection, where cameras, vibration sensors, and acoustic sensors are deployed along a production line [10]. Each sensor, empowered by an AI agent, semantically interprets its sensory stream. For instance, if one agent detects an abnormal vibration and another simultaneously identifies a visual defect, they can directly exchange semantic messages to collaboratively diagnose and localize faults in real time, without involving a remote processing center.
The I-MAC architecture provides maximum autonomy, scalability, and resilience. By minimizing reliance on centralized network infrastructure, it is particularly effective in bandwidth-constrained, delay-sensitive, or infrastructure-deficient environments. More importantly, compared to the F-MAC scheme, edge agents in I-MAC are not limited to passive data providing; they act as independent, task-aware agents with cognitive and communicative capabilities. Nonetheless, data heterogeneity and differences in prior knowledge across edge agents may lead to cognitive inconsistencies. For example, under low-light conditions, a camera-based agent may struggle to accurately perceive environmental information, whereas an infrared-based agent can operate normally. In such cases, the agents may generate divergent semantic interpretations of the same scene, potentially leading to conflicting local decisions and degraded overall system performance. Furthermore, due to hardware limitations, resource-constrained edge sensors may only support lightweight or compressed language models, which can introduce trade-offs in inference accuracy and reasoning capability.
IV-C Relay-based multimodal ISAC
Fig. 2(c) illustrates the R-MAC scheme, where each sensor is equipped with a moderately capable edge agent responsible for local feature extraction and task-relevant semantic interpretation. These edge agents extract essential information from raw multimodal data and transmit it to nearby relay nodes, such as road side units (RSU), for further processing. The relay agents then interpret, repackage, and forward the semantic information to other edge agents operating on different modality sensors. Finally, the receiving edge agent performs semantic analysis and makes task-specific decisions based on the parsed information. For example, in the drone surveillance scenario, a camera-equipped drone employs its onboard edge agent to extract positional semantics from aerial observations. This semantic information is transmitted to an RSU-based relay agent, which references a lightweight knowledge base. If the observed object is unrecognized, the relay agent formulates a semantic directive (e.g., “focus radar on coordinates (x, y)”) and forwards it to a radar sensor positioned on the ground. Upon receiving the directive, the radar’s edge agent dynamically adjusts its ISAC beam toward the specified coordinates, thus avoiding exhaustive scanning and improving sensing efficiency.
Compared to the F-MAC scheme, R-MAC supports longer distances for sensing tasks. Compared to F-MAC, it supports extended spatial coverage and reduces dependency on a central fusion node. However, the effectiveness of R-MAC depends heavily on robust SC protocols and accurate message translation between modalities. Additionally, it may underperform in scenarios that require either fine-grained, globally consistent fusion (as enabled by F-MAC) or high levels of autonomous decision-making at the edge (as in I-MAC).
To clearly illustrate the differences among the three schemes, Table I provides a comparative summary of their key characteristics. Overall, F-MAC is ideal for infrastructure-rich environments requiring centralized fusion and global decision-making (e.g., intelligent transportation hubs and smart cities). I-MAC suits infrastructure-sparse or bandwidth-limited scenarios where autonomous, peer-to-peer interaction is essential (e.g., industrial inspection sites). R-MAC bridges the two, supporting cross-modal coordination in semi-structured settings with moderate infrastructure and constrained edge devices (e.g., drone surveillance and edge patrolling systems).
TABLE I: Comparison of Three Architectural Schemes
Aspect F-MAC I-MAC R-MAC
Strengths High-accuracy inference; Strong multimodal fusion High autonomy; Strong scalability; Robust resilience Balanced workload; Moderate communication efficiency
Weaknesses Strong central dependency; Low fault tolerance Cognitive inconsistency; High edge computational burden Protocol dependency; Compromised global/local reasoning
Communication Load High Low Moderate
Fault Tolerance Low (single point of failure) High (fully decentralized) Moderate (relay-dependent)
Key Technologies LAM and SC MAS and SC SC
Use Cases Intelligent transportation hubs Industrial inspection sites Drone surveillance
V Case Study
To assess the effectiveness of the examined schemes, we conduct a case study showcasing vision-RF multimodal fusion for enhancing ISAC performance. To support efficient multimodal fusion and centralized semantic reasoning, the F-MAC scheme is adopted. As illustrated in Fig. 3(a), we consider a surveillance scenario involving a visual sensor (camera), an RF sensor (radar), a base station, and multiple sensing targets. Each sensor is equipped with a lightweight edge agent that extracts high-level features from its respective data modalities (e.g., images and radar echoes), which are subsequently transmitted to the base station via SC. At the base station, a LAM functions as the central agent, performing multimodal feature fusion and semantic-level decoding to identify the targets and estimate their motion parameters, including distance, velocity, and azimuth angle. Based on these sensing outcomes, the base station adaptively steers the ISAC beam toward the targets, thereby enhancing the overall communication rate.
Figure 3: F-MAC-based case study. (a) The illustration of the implementation scenario. (b) Key parameter settings.
V-A Experimental Settings
To construct a specialized dataset for training and evaluation, we use the VIRAT Video Dataset [11], which provides real-world surveillance video in diverse outdoor scenes. We first select three representative video clips, sampling one frame per second, resulting in approximately 10,000 RGB frames. We assume that all sensing targets in the scenes are vehicles.
To extract training labels, we first apply the YOLOv10 model [12] to identify bounding boxes of all vehicles in each frame. Next, using the Segment Anything Model (SAM) [13], we extract precise foreground vehicle segments based on those bounding boxes, generating around 800,000 labeled vehicle images for image reconstruction tasks. In addition to visual data, radar-related annotations are synthesized. We assume a fixed radar location at the lower-right corner of each image. For each vehicle, we calculate its distance and azimuth based on the relative position in the frame and assign it a velocity profile. Simulated RF signals are then generated accordingly. Some key parameter settings are shown in Fig. 3(b).
We compare it against a single-modality baseline, RF-ISAC, a conventional unimodal approach that relies exclusively on RF signals for sensing. The evaluation employs root mean squared error (RMSE) as the metric for sensing accuracy and communication rate as the metric for communication performance.
V-B Evaluation Results
Figure 4: RMSE Comparison Between F-MAC and RF-ISAC Across SNRs.
Figure 5: The communication rate comparisons between F-MAC and RF-ISAC across different SNRs.
Fig. 4 presents a quantitative comparison of the RMSE performance for motion parameter estimation between F-MAC and RF-ISAC across varying SNR levels. The results demonstrate that F-MAC consistently outperforms RF-ISAC across all three key metrics: azimuth, distance, and velocity. Specifically, F-MAC maintains remarkably stable and low RMSE values, with azimuth errors around $2.2\times 10^{-3}$, distance errors near $1.3\times 10^{-3}$, and velocity errors close to $1.2\times 10^{-3}$, while exhibiting minimal fluctuation across different SNRs. In contrast, RF-ISAC shows higher RMSE, particularly under the velocity estimation. On average and across all SNR levels, F-MAC achieves a performance gain of over 80% in RMSE reduction relative to RF-ISAC, highlighting its robustness and superior estimation capability in noisy environments. As shown in Fig. 5, it is evident that F-MAC consistently achieves a higher communication rate, approximately 304.8 Mbps, regardless of the SNR, whereas RF-ISAC maintains a rate of around 301.2 Mbps. The enhanced communication performance can be attributed to the incorporation of vision-RF multimodal fusion, which enables F-MAC to more accurately locate targets, identify their motion state, and perform more accurate ISAC beam adjustments.
These results validate the effectiveness of multimodal fusion in F-MAC. On one hand, the inclusion of the vision modality enables F-MAC to leverage richer information for sensing tasks, thereby improving sensing accuracy and robustness compared to single-modality RF-based approaches. On the other hand, based on the accurate sensing results, the base station could perform more efficient ISAC beam adjustment, thus improving the communication rate.
VI Open Issues
Despite the promising capabilities of the multimodal ISAC, several challenges remain open for future research and development.
VI-A Adaptive Multimodal Sensing
In dynamic environments, it is essential to adaptively determine which modalities should be activated at each moment or for each task. This requires intelligent policies capable of balancing sensing performance, power consumption, and network load. Therefore, designing lightweight and context-aware modality selection algorithms, possibly informed by reinforcement learning or probabilistic reasoning, remains an open challenge.
VI-B Dynamic Synesthesia Access
Synesthesia access refers to the ability of agents to selectively access different types of sensing sources depending on the situation. It is important that how to make intelligent and timely decisions about which sensors to access for a given task, taking into account the current environment, task requirements, latency constraints, and the availability of sensing resources. Hence, it is crucial to develop unified frameworks capable of dynamically orchestrating such sensor access decisions across multiple agents.
VI-C Online Learning
To maintain performance in changing environments, the agents deployed on the sensors must be capable of evolving. This accounts for efficient online learning strategies that enable the continual update of AI models with minimal communication and computation overhead. Detailed challenges include how to avoid catastrophic forgetting, preserve stability-plasticity balance, and reduce training latency in edge and collaborative settings.
VI-D Privacy and Security
Ensuring data privacy and security in sensors is a critical concern, particularly when the multimodal data includes sensitive visual, audio, or location information. Future research should explore privacy-preserving techniques, such as federated learning, differential privacy, and homomorphic encryption, adapted for multimodal and distributed architectures. Moreover, robust authentication and trust mechanisms are needed to prevent adversarial access or manipulation of sensor data and model outputs.
VII Conclusion
In this article, we first highlighted the necessity of transitioning from conventional single-modal ISAC to multimodal ISAC, driven by the growing demands of complex and dynamic 6G applications. Then, we analyzed the key challenges in enabling multimodal ISAC, including the fusion of multimodal data, the high communication overhead among distributed sensors, and the design of efficient and scalable system architectures. To address these challenges, we introduced several advanced enabling technologies, namely, LAM, SC, and MAS, which collectively provide the foundations for achieving multimodal ISAC schemes. Next, to operationalize these technologies, we zoomed into three architectural paradigms: F-MAC, I-MAC, and R-MAC, each designed to support efficient device coordination and modality collaboration under different deployment scenarios. A case study based on the F-MAC framework demonstrated substantial improvements in sensing accuracy and image reconstruction quality, achieving up to an 80% performance gain compared to conventional RF-ISAC systems. Finally, we outlined several open research challenges that must be addressed in the future.